An aqueous calcium chloride solution is 5.75% calcium chloride by mass. Calculate the mole fraction of calcium chloride in the solution.
Answer #1

Suppose weight of calcium chloride solution = 100 g then

5.75 % calcium chloride = 5.75 g of Calcium Chloride

Weight of water = (100 - 5.75) g = 94.25 g

Number of moles = mass in gram / Molar mass

Molar mass of calcium chloride = 110.98 g/mol

Molar mass of water = 18 g/mol

Number of moles of CaCl2 = 5.75 g / 110.98 g/mol = 0.05181 mol

Number of moles of water = 94.25 g / 18 g/mol = 5.236 mol

Total number of moles = (5.236 + 0.05181) mol = 5.288 mol

Mole fraction of CaCl2 = Mole of CaCl2 / Total number of moles = 0.05181 mol / 5.288 mol = 0.0098

Mole fraction of calcium chloride = 0.0098